The sum of $127 was didvided among 3 people so that the second received $5 less than twice as much as the first, and the third received $2 more than the second. How much did each person receive?
Answer by stanbon(75887)

The sum of $127 was didvided among 3 people so that the second received $5 less than twice as much as the first, and the third received $2 more than the second. How much did each person receive?
---------
Let the amount the 1st received be x dollars.
Then 2nd person received 2x-5 dollars.
And 3rd person received 2x-5+2 = 2x-3 dollars.
-------------------
EQuATION:
Total dollars was 127
x+2x-5+2x-3 = 127
5x-8=127
5x=135
x=27 dollars for 1st person
2x-5=2*27-5=49 dollars for 2nd person
2x-3=2*27-3= 51 dollars
Cheers,
Stan H.